A circular pool has a radius of 4 m and a height of 2.5 m.The pool is filled up to 3/4 of its capacity.Calculate the number of litres of water that are in the pool.
1
Expert's answer
2019-04-01T12:25:46-0400

Given that the pool has a cylindrical shape, we will find the full volume of the pool using the formula


"V=\pi*R^2*H"

where R is cylinder radius; H is cylinder height. According to the condition of the problem R=4m and H=2.5m. Therefore


"V=\pi*4^2*2.5=125.6637m^3"

The pool is filled up to 3/4 of its capacity. Hence 


"V'=3\/4*V=94.2478m^3"

Convert the resulting volume to the liters get


"V'=94247.8 literes"

Answer. 94247.8literes